# Does GPT Really Get It? A Hierarchical Scale to Quantify Human and AI’s Understanding of Algorithms

### v1

@@ Abstract
As Large Language Models (LLMs) perform (and sometimes excel at) more and more complex cognitive tasks, a natural question is whether AI really understands. The study of understanding in LLMs is in its infancy, and the community has yet to incorporate well-trodden research in philosophy, psychology, and education. We initiate this, specifically focusing on understanding algorithms, and propose a hierarchy of levels of understanding. We use the hierarchy to design and conduct a study with human subjects (undergraduate and graduate students) as well as large language models (generations of GPT), revealing interesting similarities and differences. We expect that our rigorous criteria will be useful to keep track of AI’s progress in such cognitive domains.
1 Introduction
@@
Since the release of GPT-4, mainstream users have begun to rely on Large Language Models (LLMs) for increasingly complex tasks. However, the degree to which it is safe, legal, and ethical to rely on LLMs has been under fierce debate. Across many studies, researchers have identified skills at which LLMs appear to fall short, such hallucinations, failure to plan, and lack of understanding (Rawte, Sheth, and Das 2023; Mahowald et al. 2024; Valmeekam et al. 2023). However, the literature notably lacks rigorous criteria to measure the progress toward solving these issues. A particular problem lies in claims surrounding understanding; AI understanding is frequently compared to human understanding, and it is folklore among AI researchers that the reasoning process of LLMs differ from humans. However, while the concept of understanding is widely discussed, it remains ill-defined.
@@
In this paper, we propose an algorithmic definition of understanding an algorithm with the following properties: (a) it provides a scale by which to evaluate any agent’s understanding of an algorithm, (b) aligns with the standard usage of the term ‘understanding’ in philosophy and psychology, and (c) can evaluate LLM’s progress toward understanding algorithms.
Motivation - Why study algorithm understanding?
@@
Large language models are increasingly trusted for coding assistance. Code generation tools such as GitHub Copilot (GitHub 2024) and Meta’s Code Llama (Roziere et al. 2023) are currently used in practice to improve developer productivity(Vaithilingam, Zhang, and Glassman 2022; Mozannar et al. 2024) and assist novice programmers in learning(Kazemitabaar et al. 2023; Becker et al. 2023). It is likely that the degree of AI involvement in software development will only grow as these tools improve. However, reliance on imperfect systems comes with risk. Tools such as Copilot are known to generate code that is subject to license(Becker et al. 2023) or contains security vulnerabilities(Pearce et al. 2022).
We argue that the question of whether LLMs demonstrate meaningful understanding of algorithms is relevant if we are relying on LLMs implement algorithms in production or teach them to novice programmers.
@@
We additionally argue that algorithm understanding is a distinct concept from language understanding that deserves its own line of study. Those who argue that LLMs do not understand language draw a distinction between linguistic form and meaning(Bender and Koller 2020; Mitchell and Krakauer 2023; Pavlick 2023). When humans understand language, their understanding is informed by their communicative intent and the real-life properties of the objects described. Thus, a system trained only to replicate statistical correlations between words cannot understand language in the way that humans do. Algorithms, however, can be precisely represented using formal programming languages. One could argue that a computer can meaningfully observe an algorithm in full through code implementations and examples.
That is not to say that a human’s understanding of algorithms is not informed by real-world experience. For example, a student may understand shortest-path algorithms partially through their experience using navigational apps to get to work. Nonetheless, it is easier to work around this discrepancy between humans and LLMs when the object of study is formally defined.
@@
The past few years have seen an explosion of studies exploring the ability of LLMs to answer complex mathematical questions. Researchers have developed prompting strategies to enable multi-step reasoning (Wei et al. 2022; Fu et al. 2022). Still others fine-tune models to improve mathematical provlem-solving (Yu et al. 2023; Luo et al. 2023). The benchmarks for these methods typically include large datasets such as GSM8k (Cobbe et al. 2021) (grade school word problems) and MATH (Hendrycks et al. 2021) (math competition problems). These works focus on correct evaluation and do not address whether the language models understand mathematical reasoning.
@@
Also relevant to this work is the distinction between deep and surface-level learning from educational psychology (Marton and Säljö 1976; Beattie IV, Collins, and McInnes 1997). One influential categorization comes from Mayer (2002), who categorized student learning into cognitive processes and identified testable skills which arise with understanding. This categorization partially informed the levels we define below.
2 An Algorithmic Definition
We ask the question: how well does an agent understand an algorithm? Our goal is to create a definition of understanding that is testable and algorithmic. Therefore, we adopt a functional lens, meaning that we define understanding by what it allows the agent to do.
@@ 2 An Algorithmic Definition
In this work, we ask whether an agent $\mathcal{E}$ understands a computable function $f:\Omega\rightarrow\Sigma^{*}$. By computable, we say that there exists a Turing Machine which takes $x\in\Omega$ as input and halts with $f(x)$ on its tape, using a standard definition (Sipser 1996). Let $\mathcal{A}$ be an algorithm that computes $f$.
We assume that $\mathcal{E}$ (1) has a long-term memory system and (2) can perform computation, enabled by a working memory with finite capacity $M_{\mathcal{E}}$. We say that $\mathcal{E}$ knows a function $f$ or algorithm $\mathcal{A}$ if it has a representation $R_{f}/R_{\mathcal{A}}$ stored in its long-term memory. When $\mathcal{E}$ computes the function $f$ on an input $x\in\Omega$, it runs an internal algorithm $\mathcal{A}_{\mathcal{E}}$, which may or may not be the same as $\mathcal{A}$. The agent’s understanding of $f$ or $\mathcal{A}$ will be measured by its ability to manipulate this representation to produce answers to queries. This definition is based on the Understanding as Representation Manipulability system (URM) proposed by Wilkenfield (Wilkenfeld 2013). This system and the form of the representation is discussed further in the next section.
@@
When executed by a computer, an algorithm is represented as a string of bytes that are interpreted as a series of instructions. These instructions are stored in the computer’s memory, and in this way we can say that the computer has a representation of the algorithm.
Both humans and language models also have representations of the algorithm stored in their memory. For language models, this representation is collected from the thousands of examples, explanations, and code snippets that appear in its training data. The mechanism behind human memory is not understood as precisely. However, humans also learn via hearing explanations, collecting examples, and reinforcing their knowledge. This forms a representation encoded in the neural pathways of our brains (Durstewitz, Seamans, and Sejnowski 2000; György Buzsáki 2019)
We employ the framework of Understanding as Representation Manipulability (URM)(Wilkenfeld 2013). This theory posits that understanding arises from the ability to modify the internal representation of a concept in order to make effective inferences. In our example above, while a computer possesses a representation of the algorithm, it cannot independently modify it to answer questions. On the other hand, while a human may make more mathematical mistakes, they can answer interesting questions by ‘turning the problem around’ in their head.
Wilkenfeld declines to characterize the structure of the structure of the representations, other than to state that they are “computational structures with content that are susceptible to mental transformations”(Wilkenfeld 2013). Because of this, it is a general system that is applicable to many different types of agents.
It is interesting to think of a language model in this framework because it is clear what the representation and modification should be. The representation is a vector of weights, and a question related to the algorithm is also encoded as weights. The LLM makes inferences by performing mathematical operations on this vector (in essence, manipulating it).
@@
In this section, we define understanding as a spectrum by presenting a series of levels. Understanding at each level is loosely more difficult than the previous one, although they do not formally follow each other. Rather, they measure increasing levels of abstraction. To demonstrate the ideas, we also provide examples of questions which could be successfully answered by an agent that understands the Euclidean algorithm at each level.
For simplicity, we present these levels as deterministic; however, they can be defined with a failure probability dependent on the agent’s internal randomness and the required memory and time.
At the first level (denoted Level 1), the agent is capable of evaluating the algorithm on some ‘simple’ examples. At this level, the agent has some representation of the input-output mapping, whether or not it can formally express it.
@@
At the next level, the agent can describe how it evaluates $f(x)$ in a language that it knows. Level 2 requires the agent to output the execution steps of the algorithm on $x$ as well as produce the correct answer.
@@
$\mathcal{E}$ understands $\mathcal{A}$ at Level 2 if, given an $x\in\Omega$ with $M(x)\leq M_{\mathcal{E}}$:
  • a list of instructions
  • a description in structured natural language
  • a flow chart or other unambiguous pictoral representation of the instructions
that are run when executing $\mathcal{A}$ on $x$.
@@
The next level will take this one step further, requiring the agent to produce a set of instructions that can be followed to produce the right answer for any $x\in\Omega$.
@@
The first three levels measure the ability of the agent to recall a procedure and execute a known set of instructions. We place these in the category of ‘shallow learning’; in Mayer’s taxonomy, they fall under the cognitive processes of recognizing, recalling, and executing (Mayer 2002). Note that all three levels could be achieved by a hard-coded script.
The next two levels target deep learning, and measure cognitive processes in the ‘Understanding’ and ‘Analyzing’ categories. We split the next levels into two subtrees, broadly corresponding to examples and explanation. This approach allows us to distinguish cognitive processes utilizing functional linguistic skills from those utilizing mathematical reasoning (Mahowald et al. 2024).
At Level 4, the agent can generate examples, convert between representations, and produce descriptions at different levels of detail.
@@
Given a property $P$ of the execution path of the algorithm, $\mathcal{E}$ can generate an $x\in\Omega$ which satisfies $P$ or report that none exists. Example: Give an integer $550<x<$ 55 that requires the greatest number of recursive steps to compute GCD$(55,x)$. Describe how you chose this number.
Definition (Level 4b: Summarization).
$\mathcal{E}$ can produce descriptions of $\mathcal{A}$ satisfying the following criteria.
  • When given a portion of the code for $\mathcal{A}$ in a formal language that it knows, the agent can summarize it in natural language. Similarly, it can produce correct, executable code when given a summary of a portion of the algorithm.
  • The agent can produce natural language summaries of the algorithm at distinct levels of abstraction. A more abstract summary replaces sections of the control flow with references to other algorithms $\mathcal{E}$ can describe.
@@
At Level 5, the agent can reason on perturbations of the algorithm.
@@
Given an algorithm $\mathcal{A^{\prime}}$, the agent can determine whether $\mathcal{A}$ and $\mathcal{A^{\prime}}$ produce the same output on all $x\in\Omega$. If not, it can find a counterexample such that $\mathcal{A}^{\prime}(x)\neq f(x)$.
Example: Suppose GCD$(a,b)=x$ and GCD$(b,c)=y$. Does GCD$(a,b,c)$ = GCD$(x,y)$? Explain your reasoning.
@@
Given an algorithm $\mathcal{A^{\prime}}$, $\mathcal{E}$ can answer natural language queries about $\mathcal{A^{\prime}}$ phrased as counterfactual questions about $\mathcal{A}$.
@@
A counterfactual question asks ‘what if?’ and presents a modification to the basic assumptions of the setting. Counterfactual questions are considered in the philosophy of understanding to separate ‘knowing why’ from ‘understanding why’(Woodward 2005; Páez 2019). Páez (2019) gives the example of understanding climate change. A person might know that global warming is caused by greenhouse gases being released into the atmosphere, and could give a well-reasoned explanation of this phenomenon when asked. However, the same person may not be able to answer questions such as “what would happen to global temperatures if all human activity were to cease?” However, a climatologist, who likely understands why global warming occurs, could reason about this alternate scenario and provide an answer.
In terms of algorithms, we interpret this as the ability to answer questions about hypothetical modifications to the procedure, the input space, or the output space. In the example given, $\mathcal{A^{\prime}}$ restricts the input space to consecutive Fibonacci numbers. The task requires the agent to reason about changes in runtime under this restricted space. These levels of understanding and the partial ordering among them is summarized in Fig. 1.
@@
3 Methods
We use two classical algorithms to test our scale of understanding: the Euclidean algorithm for computing the greatest common denominator of two integers, and the Ford-Fulkerson algorithm for computing the maximum flow between two nodes on a directed graph with capacity constraints (Ford and Fulkerson 1956). Both algorithms are widely taught in undergraduate computer science curricula. The Euclidean algorithm is relatively simple, and it is the first introduction to the concept of algorithm for many students. The Ford-Fulkerson algorithm, on the other hand, is more sophisticated and is typically introduced in a full-fledged algorithms course.

### v3

@@ Abstract
As Large Language Models (LLMs) are used for increasingly complex cognitive tasks, a natural question is whether AI really understands. The study of understanding in LLMs is in its infancy, and the community has yet to incorporate research and insights from philosophy, psychology, and education. Here we focus on understanding algorithms, and propose a hierarchy of levels of understanding. We validate the hierarchy using a study with human subjects (undergraduate and graduate students). Following this, we apply the hierarchy to large language models (generations of GPT), revealing interesting similarities and differences with humans. We expect that our rigorous criteria for algorithm understanding will help monitor and quantify AI’s progress in such cognitive domains.
Introduction
@@
Since the release of GPT-4, mainstream users have begun to experiment with Large Language Models (LLMs) on increasingly complex tasks. However, the degree to which it is safe, legal, and ethical to rely on LLMs has been under fierce debate. Across many studies, researchers have identified apparent shortcomings of LLMs including hallucinations, inability to plan, and lack of understanding (Rawte, Sheth, and Das 2023; Mahowald et al. 2024; Valmeekam et al. 2023). However, the literature notably lacks rigorous criteria to measure the progress toward solving these issues. A particular problem lies in claims surrounding understanding; AI understanding is frequently compared to human understanding, and it is folklore among AI researchers that the reasoning processes of LLMs differ from those of humans. While the concept of understanding is widely discussed, it remains ill-defined.
@@
In this paper, we propose an precise definition of understanding an algorithm with the following properties: (a) it provides a scale by which to evaluate any entity’s understanding of an algorithm, (b) it aligns with the standard usage of the term ‘understanding’ in philosophy and psychology, and (c) it can be used to evaluate AI’s progress toward understanding algorithms.
Motivation: Why Study Algorithm Understanding?
@@
Large language models are increasingly trusted for coding assistance. Code generation tools such as GitHub Copilot (GitHub 2024) and Meta’s Code Llama (Roziere et al. 2023) are currently used in practice to improve developer productivity (Vaithilingam, Zhang, and Glassman 2022; Mozannar et al. 2024) and assist novice programmers in learning (Kazemitabaar et al. 2023; Becker et al. 2023). It is likely that the degree of AI involvement in software development will only grow as these tools improve. However, reliance on imperfect systems comes with risk. Tools such as Copilot are known to generate code that is subject to license (Becker et al. 2023) or contains security vulnerabilities (Pearce et al. 2022). The question of whether LLMs demonstrate meaningful understanding of algorithms is relevant if we are relying on LLMs to implement algorithms in production or teach them to novice programmers.
@@
Algorithm understanding is distinct from language understanding and deserves its own line of study. Those who argue that LLMs do not understand language draw a distinction between linguistic form and meaning(Bender and Koller 2020; Mitchell and Krakauer 2023; Pavlick 2023). When humans understand language, their understanding is informed by their communicative intent and the real-life properties of the objects described. Thus, a system trained only to replicate statistical correlations between words cannot understand language in the way that humans do. Algorithms, however, can be precisely represented using formal programming languages. One might argue that a computer can meaningfully observe an algorithm in full through code implementations and examples.
@@
The past few years have seen an explosion of studies exploring the ability of LLMs to answer complex mathematical questions. Researchers have developed prompting strategies to enable multi-step reasoning (Wei et al. 2022; Fu et al. 2022). Still others fine-tune models to improve mathematical problem-solving (Yu et al. 2023; Luo et al. 2023). The benchmarks for these methods typically include large datasets such as GSM8k (Cobbe et al. 2021) (grade school word problems) and MATH (Hendrycks et al. 2021) (math competition problems). These works focus on correct evaluation and do not address whether the language models understand mathematical reasoning.
@@
Also relevant to this work is the distinction between deep and surface-level learning from educational psychology (Marton and Säljö 1976; Beattie, Collins, and McInnes 1997). Perhaps the most influential categorization of educational goals is Bloom’s Taxonomy (Bloom et al. 1956). This taxonomy has been revisited many times since its publication; notably, Mayer (2002) categorized student learning into cognitive processes and identified testable skills which arise with understanding.
A Definition of Understanding
We ask the question: how well does an entity understand an algorithm? Our goal is a definition of understanding that is itself algorithmically testable. Therefore, we adopt a functional lens, meaning that we define understanding by what it allows the entity to do.
@@ A Definition of Understanding
In this work, we ask whether an entity $\mathcal{E}$ understands a computable function $f:\Omega\rightarrow\Sigma^{*}$. By computable, we say that there exists a Turing Machine which takes $x\in\Omega$ as input and halts with $f(x)$ on its tape, using a standard definition (Sipser 1996). Let $\mathcal{A}$ be an algorithm that computes $f$.
We assume that an entity $\mathcal{E}$ (a) has a long-term memory system and (b) can perform computation, enabled by a working memory with finite capacity $M_{\mathcal{E}}$. We say that $\mathcal{E}$ knows a function $f$ or algorithm $\mathcal{A}$ if it has a representation $R_{f}/R_{\mathcal{A}}$ stored in its long-term memory. When $\mathcal{E}$ computes the function $f$ on an input $x\in\Omega$, it runs an internal algorithm $\mathcal{A}_{\mathcal{E}}$, which may or may not be the same as $\mathcal{A}$. The entity’s understanding of $\mathcal{A}$ will be measured by its ability to manipulate this representation to produce answers to queries. This definition is based on the Understanding as Representation Manipulability system (URM) proposed by Wilkenfield (Wilkenfeld 2013).
For a particular input $x$, let $M(x)$/$T(x)$ be the memory/time required to evaluate the algorithm on $x$.
The execution path of $\mathcal{A}$ on an input $x$ is the sequence of states taken by the algorithm when executing on $x$. The trace of the execution is the execution path plus the contents of the tape at each step. Finally, define a property to be a function mapping the trace or execution path to $\{0,1\}$.
@@
For our hierarchy, we employ the framework of Understanding as Representation Manipulability (URM) (Wilkenfeld 2013). This theory posits that understanding arises from the ability to modify the internal representation of a concept in order to make effective inferences.
For language models, the representation of a concept (such as an algorithm) is collected from the thousands of examples, explanations, and code snippets that appear in its training data. The mechanism behind human memory is not understood as precisely. However, humans also learn via hearing explanations, collecting examples, and reinforcing their knowledge. This forms a representation encoded in the neural pathways of our brains (Durstewitz, Seamans, and Sejnowski 2000; Buzsáki 2019). The goal of our hierarchy will be to test how the existing internal representation can be manipulated to produce responses at different levels of difficulty.
In his description of URM, Wilkenfeld declines to characterize the structure of the representations, other than to state that they are “computational structures with content that are susceptible to mental transformations”(Wilkenfeld 2013). Thus, the definition of understanding is independent of the entity’s internal structure and the mechanism for inference. This is in line with our functional lens; the level of understanding is based on the entity’s ability to manipulate its representation to perform tasks at different levels of difficulty.
@@
In this section, we define understanding as a spectrum by presenting a series of levels. Understanding at each level is intended to be more difficult than the previous one, although they do not formally follow each other. Rather, they measure increasing levels of abstraction. To demonstrate the ideas, we also provide examples of questions that would be successfully answered by an entity that understands the Euclidean algorithm for GCD at each level. The full hierarchy is summarized in Figure 1.
For simplicity, we present these levels as deterministic; however, they can be defined with a failure probability dependent on the entity’s internal randomness and the required memory and time.
At the first level (denoted Level 1), the entity is capable of evaluating the algorithm on some ‘simple’ examples, where the simplicity of an input is defined by the length of the execution path. At this level, the entity has some representation of the input-output mapping, whether or not it can formally express it.
@@
At the next level, the entity can describe how it evaluates $f(x)$ in a language that it knows. Level 2 requires the entity to output the execution steps of the algorithm on $x$ as well as produce the correct answer.
@@
$\mathcal{E}$ understands $\mathcal{A}$ at Level 2 if, given an $x\in\Omega$ with $M(x)\leq M_{\mathcal{E}}$ it can provide one of the following:
  • the execution path in natural language or code
  • a flow chart or other unambiguous pictoral representation of the execution path
executed when running $\mathcal{A}$ on $x$.
@@
The next level will take this one step further, requiring the entity to produce a set of instructions that can be followed to produce the right answer for any input $x\in\Omega$.
@@
The first three levels measure the ability of the entity to recall a procedure and execute a known set of instructions. We place these in the category of ‘shallow learning’; in Mayer’s taxonomy, they fall under the cognitive processes of recognizing, recalling, and executing (Mayer 2002). Note that all three levels could be achieved by a hard-coded script.
The next two levels target deep learning, and measure cognitive processes in the ‘Understanding’ and ‘Analyzing’ categories. We split the next levels into two subtrees, to distinguish cognitive processes utilizing functional linguistic skills from those utilizing mathematical reasoning (Mahowald et al. 2024).
At Level 4, the entity demonstrates an understanding of ‘why’ the algorithm is constructed as it is. It requires them to provide an example to illustrate a property (mathematical reasoning) or explain the existence of a property to a specified audience (reasoning with language).
@@
Given a property $P$ of an execution path of the algorithm $\mathcal{A}$, $\mathcal{E}$ can generate an $x\in\Omega$ which satisfies $P$ or report that none exists. Example: Give an integer $550<x<$ 55 that requires the greatest number of recursive steps to compute GCD$(55,x)$. Describe how you chose this number.
Definition (Level 4b: Explanation).
Given $\mathcal{A}$, or a property $P$ satisfied by the execution path of $\mathcal{A}(x)$, and an audience $\mathcal{E}^{\prime}$, the entity can produce a text in natural language that has the following characteristics:
  • Accurately describes the steps of the algorithm/execution path.
  • Abstracts or shortens the full description by referencing other algorithms known by the audience.
  • Uses examples and analogies to other algorithms known by the audience to convey intuition.
@@
At Level 5, the entity can reason on perturbations of the algorithm and perturbations of the input, and it can describe the effect on the execution path. Under reasoning with mathematics, this includes skills such as certifying if a modification to an algorithm changes the output for a subset of examples. Under reasoning with language, this includes describing the effects of modifying inputs, or answering counterfactual questions about modifications to the algorithm.
@@
The entity can answer questions about $\mathcal{A}$ of the following form.
  • Given an algorithm $\mathcal{A^{\prime}}$, the entity can determine whether $\mathcal{A}$ and $\mathcal{A^{\prime}}$ produce the same output on all $x\in\Omega$. If not, it can find a counterexample such that $\mathcal{A}^{\prime}(x)\neq\mathcal{A}(x)$.
  • Given a relation $R\subset\Omega\times\Omega$, the entity can find a pair $(x,x^{\prime})\in R$ with different execution paths on $\mathcal{A}$.
Example: Determine whether the following statement is true. If not, provide a counterexample. If $x>y$, then computing $GCD(2x,y)$ with the Euclidean algorithm requires more division operations than computing $GCD(x,y)$.
@@
The entity can produce natural language descriptions of $\mathcal{A}$ of the following form.
  • Given an algorithm $\mathcal{A^{\prime}}$ and an audience $\mathcal{E}^{\prime}$, the entity can produce an explanation (c.f. Level 4b) contrasting the two algorithms.
  • Given a relation $R\subset\Omega\times\Omega$, the entity can describe a property highlighting the differences in execution paths for $(x,x^{\prime})\in R$.
@@
Hypotheses
We conducted an experiment on LLMs and human participants with two main goals; 1) to assess the proposed hierarchical scale (Figure 1) as a tool for comparing levels of understanding, and 2) to rate algorithm understanding across generations of GPT. We will assess the scale with a student survey, where we can use educational level as a basis for comparison. Then, we will apply the same questions to GPT and assess its understanding on the same scale. Related to these goals, we test the following hypotheses:
We expect the fraction of correct answers to be non-increasing with higher levels of understanding. Furthermore, more education and training in algorithms should be reflected in the scores, so we expect graduate students to perform better than undergraduates.
2. Newer generations of GPT understand algorithms at a higher level than older generations.
Concretely, we expect an increase in performance at higher levels between GPT-3.5 and GPT-4.
3. LLMs will exhibit a performance gap between natural language reasoning and mathematical reasoning tasks.
We expect the difference in performance between these two types of tasks to be much smaller in students than LLMs. Furthermore, we expect that GPT may have a higher performance at Level 3, since GPT is fine-tuned on code generation and has been exposed to code for common algorithms.
Methods
We use two classical algorithms to test our scale of understanding: the Euclidean algorithm for computing the greatest common divisor of two integers, and the Ford-Fulkerson algorithm for computing the maximum flow between two nodes on a directed graph with capacity constraints (Ford and Fulkerson 1956). Both algorithms are widely taught in undergraduate computer science curricula. The Euclidean algorithm is relatively simple, and it is the first introduction to the concept of algorithm for many students. The Ford-Fulkerson algorithm, on the other hand, is more sophisticated and is typically introduced in an upper-level algorithms course.
Experimental Design
In this section, we describe how the assessment was constructed. This will serve as a guide to generalize the experiment to other algorithms.
For each of the assessed algorithms, we produced a series of questions corresponding to each of the levels (Figure 1).
  1. 1.
    A trivial instance of the problem. If the entity understands the input and output space, the problem can be answered without calculation.
  2. 2.
    An intermediate instance of the problem. This is answerable without a calculator for most undergraduates, but requires the entity to run some internal algorithm to compute each individual step.
  3. 3.
    A coding problem. This problem requires the entity to translate a part of the algorithm to code.
  4. 4.
    Either an example problem (a) or an explanation problem (b). This question asks the entity to provide an example to illustrate a concept or explain the algorithm to a specified audience.
  5. 5.
    Either a counter-example (a) or extension problem (b). This problem asks the entity to reason about modifications to the algorithm through calculation or explanation.
Examples of questions posed in the survey are included under the level definitions, and the full survey will be available in the Appendix. The most challenging part of constructing the assessment is in questions 4 and 5. An algorithm can be thought of as consisting of three parts - an input space, an output space, and a transformation procedure. The procedure can be further broken up into subroutines consisting of simpler algorithms that an entity may understand in other contexts. To construct an explanation problem, we specify an audience for the explanation, which cues the level of detail and the types of subroutines which can be referred to. We construct a counterfactual problem by modifying the input, output, or a key subroutine.